a sphere has a radius of 4. find the surface area
Answer by stanbon(75887)

a sphere has a radius of 4. find the surface area
------------------
Surface Area = 4*pi*r^2
Surface Area = 4*pi*16 = 64pi sq. units
========================
Cheers,
Stan H.